# Sharp thresholds for spanning regular subgraphs

### v2

@@ 1. Introduction
Since the original paper of Erdős and Rényi [8] the task of determining the asymptotic behaviour of $p_{c}(\mathcal{Q})$ for increasing properties $\mathcal{Q}$ has been a central topic in probabilistic combinatorics. While the asymptotic order of the probability threshold has been determined for many natural increasing graph properties, a general solution remains unknown, and determining the exact asymptotics of $p_{c}(\mathcal{Q})$ is even more challenging. In this paper we address the latter question, explicitly posed in [29, Question 2]. We provide an answer for a class of increasing properties generated by $d$-regular graphs and resolve a notable conjecture in this area regarding the asymptotics of $p_{c}$ for the appearance of the second power of a Hamilton cycle.
Despite this challenge, for any arbitrary increasing $\mathcal{Q}$, the value of $p_{c}(\mathcal{Q})$ can be established up to a $\log n$-factor if the so called expectation threshold is known — thanks to the remarkable result of Park and Pham [28], which resolved the renowned conjecture of Kahn and Kalai. Let us recall the definition of the expectation threshold and then state the theorem of Park and Pham. Roughly speaking, the expectation threshold $p_{e}(\mathcal{Q})$ is the moment when the expected number of graphs generating $\mathcal{Q}$ hits $1/2$. Formally, $p_{e}(\mathcal{Q})$ is the maximum $p$ such that there exists a set of graphs $\mathcal{Q}^{\prime}$ satisfying $\sum_{G\in\mathcal{Q}^{\prime}}p^{|E(G)|}\leq\frac{1}{2}$ and $\mathcal{Q}\subseteq\langle\mathcal{Q}^{\prime}\rangle$, where $\langle\mathcal{Q}^{\prime}\rangle$ is the upwards closure of $\mathcal{Q}^{\prime}$. Due to Markov’s inequality, $p_{e}(\mathcal{Q})\leq p_{c}(\mathcal{Q})$. Let $n_{\mathcal{Q}}$ be the number of edges in the largest minimal element of $\mathcal{Q}$.
@@
The upper bound in this theorem is tight: there exists an increasing $\mathcal{Q}$ such that $p_{c}(\mathcal{Q})=\Theta(p_{e}(\mathcal{Q})\cdot\log n_{\mathcal{Q}})$. For example, let $\mathcal{Q}$ be the property of being Hamiltonian. It is a routine to check that $p_{e}=\Theta(1/n)$, while $p_{c}=(1+o(1))\frac{\ln n}{n}$ [21, 30]. It is fair to say that the main remaining challenge is to classify properties based on the value of the ratio $\frac{p_{c}(\mathcal{Q})}{p_{e}(\mathcal{Q})}\geq 1$, see, e.g., [3]. In particular, for which properties $\mathcal{Q}$ is $p_{c}(\mathcal{Q})=(1+o(1))p_{e}(\mathcal{Q})$? In this paper, we explore this question for the property of containing a given (unlabelled) regular spanning subgraph.
@@
Let $d\geq 3$ be a fixed constant. Let $F=F(n)$ be a sequence of $d$-regular graphs on the vertex set $[n]:=\{1,\ldots,n\}$. For $\tilde{F}\subset F$, we call a vertex $v\in V(\tilde{F})$ boundary, if it has degree strictly less than $d$ in $\tilde{F}$. We call the vertex boundary of $\tilde{F}$ the set $\partial_{v}(\tilde{F})$ of all boundary vertices and the edge boundary the set $\partial_{e}(\tilde{F})$ of all edges of $F$ between $\partial_{v}(\tilde{F})$ and $V(F)\setminus V(\tilde{F})$. Let the increasing property $\mathcal{Q}_{F}$ be generated by the family $\mathcal{F}_{n}$ of all isomorphic copies of $F$ on $[n]$. For simplicity, we denote $p_{c}(F):=p_{c}(\mathcal{Q}_{F})$ and $p_{e}(F):=p_{e}(\mathcal{Q}_{F})$. We also let $X_{F}$ be the number of graphs from $\mathcal{F}_{n}$ that are subgraphs of $\mathbf{G}\sim G(n,p)$. Since
@@
we get that $p_{c}(F)\geq p_{e}(F)\geq(1-o(1))(e/n)^{2/d}$. On the other hand, if every subgraph of $F$ has edge boundary of size at least $d$, then $p_{e}(F)=O(n^{-2/d})$ since it is not hard to see that $\mathcal{F}_{n}$ is $O(n^{-2/d})$-spread, i.e. for every graph $H\subset F$,
@@
if $F$ satisfies a stronger condition on the edge boundary: every $\tilde{F}\subset F$ with $3\leq|V(\tilde{F})|=o(\log n)$ satisfies $|\partial_{e}(\tilde{F})|\geq 2d$. In particular, for powers of a Hamilton cycle [^3] For a graph $G$, its $k$-th power $G^{k}$ is obtained by adding to $G$ edges between vertices that are at distance at most $k$ in the graph metric induced by $G$., this result implies the following: for every $k\geq 3$, the threshold probability for containing the $k$-th power of a Hamilton cycle equals $\Theta(n^{-1/k})$. However, the proof of Riordan does not work for $k=2$ since the minimum size of the edge boundary of a non-trivial subgraph of the second power of a Hamilton cycle equals $d+2=6<2d$. We also note that, for a $d$-regular graph to satisfy the condition of Riordan, $d$ should be at least 4. In [23], Kühn and Osthus proved that $n^{-1/2+o(1)}$ is the threshold probability for containing the second power of a Hamilton cycle and conjectured that the threshold is actually $\Theta(n^{-1/2})$. In [27], Nenadov and Škorić proved the upper bound $n^{-1/2}(\log n)^{4}$, which was improved to $n^{-1/2}(\log n)^{3}$ by Fischer, Škorić, Steger, and Trujić in [11], and to $n^{-1/2}(\log n)^{2}$ in an unpublished work of Montgomery (see [16]). From Theorem 1.1 (as well as from its weaker fraction version [12]), it follows that the threshold is $O(n^{-1/2}\log n)$. The conjecture of Kühn and Osthus was eventually resolved by Kahn, Narayanan, and Park in 2020 [18]. They further conjectured that $p_{c}(F)=(1+o(1))\sqrt{e/n}$. This paper resolves the conjecture, see Section 1.3.
On the other hand, if $F$ has many small subgraphs with smaller edge boundaries, then it may happen that $p_{c}(F)\gg p_{e}(F)$. For instance, assume that there exists $v=v(n)=o(\log n)$ and a graph $H=H(n)$ on $[v]$ with $\frac{dv}{2}-\frac{d}{2}$ edges such that every vertex of $F$ belongs to a subgraph which is isomorphic to $H$ (clearly, its edge boundary equals $d$). Then, an increasing (at most logarithmic) factor arises — for constant $v$ see the proof in [32] which can be directly generalised to all $v=o(\log n)$. Actually, if we allow even smaller edge boundaries, then the threshold probability may increase by a power of $n$, as well as the expectation threshold.
@@
As follows from the above discussion, our bound on edge boundaries is tight, that is, in some sense, the condition on the edge boundary is the only obstacle in getting the equality (1). Indeed, for every $3\leq v=o(\log n)$ and every $0\leq\ell\leq d$ (such that $dv-\ell$ is even), there exists a graph $F$ that have a subgraph on $v$ vertices with edge boundary of size $\ell$ such that (1) does not hold. The proof of Theorem 1.2 uses similar techniques as in [7, 18] and does not require accurate spread approximations, for this reason we postpone it to Appendix A.
@@
Let $\mathbf{G}\sim G(n,p)$. The threshold probability $p_{c}(\mathcal{Q})$ is called sharp, if, for every constant $\varepsilon>0$, the following is true: if $p>(1+\varepsilon)p_{c}(\mathcal{Q})$, then whp $\mathbf{G}\in\mathcal{Q}$; if $p<(1-\varepsilon)p_{c}(\mathcal{Q})$, then whp $\mathbf{G}\notin\mathcal{Q}$. Friedgut proved [14] that all increasing (and isomorphism-closed) properties that do not have sharp thresholds are essentially determined by the presence of subgraphs of bounded sizes. Omitting technical details, it means that, for some $p=\Theta(p_{c}(\mathcal{Q}))$, $\mathbf{G}\sim G(n,p)$ sprinkled with $\mathbf{G}^{\prime}\sim G(n,\varepsilon p)$, for some $\varepsilon>0$, is less likely to have the property $\mathcal{Q}$, than the union of $\mathbf{G}$ with an independently sampled random clique of some bounded size (see, e.g., [13, Theorem 2.3]). With this remarkable result, it is fairly straightforward to check that the threshold probabilities in Theorem 1.2 are sharp. To the best of our knowledge, there are no general results describing asymptotics of sharp $p_{c}$ for a wide class of increasing properties (say, for $\mathcal{Q}_{F}$, where $F$ are $d$-regular). In this paper, we find asymptotics of $p_{c}(F)$ for most $d$-regular graphs $F$.
@@
Note that, for every $d\geq 4$, almost all $d$-regular graphs satisfy the first condition in Theorem 2 (see [5, 19, 24]). Indeed, for every $d\geq 3$, letting $\mathbf{G}_{d}$ be a random $d$-regular graph on $[n]$, whp $\mathbf{G}_{d}$ is asymmetric [24], implying 1.(a). Moreover, whp $\mathbf{G}_{d}$ has Cheeger constant $\Theta(1)$ and, whenever $d\geq 4$, whp $\mathbf{G}_{d}$ does not have subgraphs with $33\leq v\leq n-$ 3 vertices and the edge boundary of size at most $d$, implying 1.(b). If $d=3$, then whp there are no subgraphs with $33\leq v\leq n-$ 3 vertices and the edge boundary of size at most $d$ other than $K_{3}$ and their vertex-complements. Therefore, whp $\mathbf{G}_{d}$ satisfies the first condition, if it does not contain triangles. Also, $k$-th powers of Hamilton cycles for every $k\geq 3$ and toroidal square grids $T_{m,n/m}\cong C_{m}\square C_{n/m}$ for $m\geq 4$ satisfy the second condition as well. It is also easy to see that the second condition implies sharp thresholds for square and triangular lattices (by considering their regular completions). It is worth recalling that thresholds for lattices were investigated in the past: the question of which random graphs contain spanning lattices was raised by Venkatesan and Levin [36]. Thresholds for lattices, up to a polylog-factor, were proved in [10] and, up to a constant factor, were proved in [31] (they also follow from Theorem 1.2 and [7, Corollary 1.5]). Summing up, we get the following corollary that demonstrates that the set of regular graphs $F$ satisfying the restrictions in Theorem 2 is quite broad.
@@
Examples of graphs that do not satisfy restrictions from Theorem 2 include the second power of a cycle, toroidal grid $T_{3,n/3}$, and, so-called, 2-overlapping 4-cycles. The latter graph, denoted by $C^{e}_{4,n}$, is obtained from $n/2$ cyclically order copies of $C_{4}$, where two consecutive $C_{4}$ overlap in exactly one edge, whereby each $C_{4}$ overlaps with two copies of $C_{4}$ in opposite edges. Recall that Kahn, Narayanan, and Park in [18] proved (1) for the second power of a cycle $F$, resolving the conjecture of Kühn and Osthus [23]. Moreover, using the same method, Espuny Díaz and Person in [9] proved (1) for $F=C^{e}_{4,n}$ answering the question of Frieze [15] (we note that (1) for these graphs also follow both from Theorem 1.2 and [7, Corollary 1.5]).
1.3. Second powers of Hamilton cycles.
@@
In 2020, Kahn, Narayanan, and Park [18] conjectured that (2) holds when $F$ is the second power of a cycle. Due to the natural barrier in the application of the fragmentation technique that was used to prove (1) for the second power of a Hamilton cycle [18], to resolve the conjecture of Kahn and Kalai [28], and to prove the so called spread lemma [2] (see details in Section 1.4), this conjecture attracted significant interest and was reiterated in [16, 29]. In this paper, we prove the conjecture.
@@
Let $F=F(n)$ be the second power of an $n$-cycle on $[n]$. Then $F$ satisfies (2).
@@
Let $d\geq 3$ be a constant, $r=r(n)=o(\log\log n)$ [^4] We did not try to optimise the bound on $r$ and we believe that it can be improved using our methods., and $F=F(n)$ be a sequence of graphs on $[n]$ satisfying the following.
@@
In particular, (2) holds for the two other ‘naive’ cases that Theorem 2 does not cover — $F=C^{e}_{4,n}$ and $F=T_{3,n/3}$.
@@
In order to prove Theorem 1.7, we introduce a novel technique of choosing fragments — we show that the “cyclic” structure of second powers of Hamilton cycles allows to choose rare fragments that are sparser than typical ones, from sufficiently many copies of $F$. The key idea from [18] that proves (1) for second powers of cycles is as follows. Fix any $F\in\mathcal{F}_{n}$, before the edges of the uniformly random graph $\mathbf{W}\sim G(n,m)$ are exposed, where $m=\lfloor p{n\choose 2}\rfloor$. Consider $F\cup\mathbf{W}$ and show that whp almost every $F^{\prime}\subset F\cup\mathbf{W}$ from $\mathcal{F}_{n}$ is such that $|F^{\prime}\setminus\mathbf{W}|=O(\sqrt{n})$. It means that we may replace almost every $F\in\mathcal{F}_{n}$ with its fragment $F^{\prime}\setminus\mathbf{W}\subset F$ — for typical $F^{\prime}\subset F\cup\mathbf{W}$ — of size $O(\sqrt{n})$ and expose another $\mathbf{\tilde{W}}\sim G(n,\tilde{m})$ independently where $\tilde{m}$ is large enough to cover at least one fragment whp. Then whp $\mathbf{W}\cup\mathbf{\tilde{W}}$ contains a graph from $\mathcal{F}_{n}$.
However, if a fragment contains a subgraph with edge boundary of size $d+2=6$ with $\Omega(\ln n)$ edges — such a subgraph is exactly the second power of a path, we call it closed — then $\tilde{m}=\lceil\delta n\sqrt{n}\rceil$ additional uniformly random edges with a certain bounded from zero $\delta>0$ is not enough to cover it. Unfortunately, we suspect that typical fragments do contain closed subgraphs of size $\Omega(\ln n)$. So, this approach does not allow to get a sharp bound on the threshold. This is the main complication that does not allow to apply directly the fragmentation technique to resolve the conjecture of Kahn, Narayanan, and Park. A careful implementation of the technique gives upper bound on the threshold $(\sqrt{e}+1+o(1))n^{-1/2}$, see Remark 1.3.
@@
The proof of Theorem 2 is much more straightforward — typical fragments are good enough to derive (2). Indeed, both conditions do not permit closed subgraph of size $\Omega(\log n)$. We prove the first part of Theorem 2 via $1/\delta$ fragmentation steps. The proof of the second part of Theorem 2 requires only two fragmentations. Actually the usual second moment method works in this case — for the uniform model instead of the binomial, as in the paper of Riordan [31]. It is possible to show that $\mathrm{Var}X_{F}=O((\mathbb{E}X_{F})^{2})$ and then to apply the powerful results of Friedgut [13, 14]. In particular, this strategy was used by Narayanan and Schacht to determine sharp thresholds for nonlinear Hamiltonian cycles in random hypergraphs [27]. However, we give the proof of the second part of Theorem 2 using fragmentation for the sake of convenience and coherence.
@@
Let us finally say a few words about our proof of Theorem 1.2. Kahn, Narayanan, and Park in [18] noted that the key observation enabling their proof that the threshold for the appearance of the second power of a Hamilton cycle $F$ equals $p_{c}(F)=\Theta(n^{-1/2})$ is that $\mathcal{F}_{n}$ is $(1+o(1))\sqrt{e/n}$-spread. They refined the concept of spreadness by incorporating the number of components in a subgraph. This refinement was later distilled by Espuny Díaz and Person in [9] under the name of superspreadness and was used to extend the result of Kahn, Narayanan, and Park to a broader class of spanning subgraphs in $G(n,p)$. Somewhat surprisingly, this property allows to prove the weaker result [7, Corollary 1.5] but does not allow to prove Theorem 1.2, as it treats contributions from connected components of both bounded and growing sizes in the same manner. It is also worth noting that Spiro [33] proposed another generalisation of spreadness, which, in particular, recovers the results from [9] and served as a key ingredient in the proofs of [7].

### v3

@@ 1. Introduction
Since the original paper of Erdős and Rényi [9] the task of determining the asymptotic behaviour of $p_{c}(\mathcal{Q})$ for increasing properties $\mathcal{Q}$ has been a central topic in probabilistic combinatorics. While the asymptotic order of the probability threshold has been determined for many natural increasing graph properties, a general solution remains unknown, and determining the exact asymptotics of $p_{c}(\mathcal{Q})$ is even more challenging. In this paper we address the latter question, explicitly posed in [31, Question 2]. We provide an answer for a class of increasing properties generated by $d$-regular graphs and resolve a notable conjecture in this area regarding the asymptotics of $p_{c}$ for the appearance of the square of a Hamilton cycle. In particular, our results imply that
  • for (asymptotically) almost all $d$-regular graphs $F$ on $\{1,\ldots,n\}$, the threshold for containing a spanning subgraph isomorphic to $F$ is $(1+o(1))(e/n)^{2/d}$;
  • the thresholds for containing a square lattice and a triangular lattice are, respectively, $(1+o(1))(e/n)^{1/2}$ and $(1+o(1))(e/n)^{1/3}$;
  • for every integer $k\geq 2$, the threshold for containing the $k$-th power of a Hamilton cycle is $(1+o(1))(e/n)^{1/k}$, thus resolving the conjecture of Kahn, Narayanan, and Park [20].
1.1. Expectation threshold.
Although pinning down exact asymptotics for probability thresholds is notoriously difficult, for any increasing property $\mathcal{Q}$ the value of $p_{c}(\mathcal{Q})$ can be established up to a $\log n$-factor if the so called expectation threshold is known — thanks to the remarkable result of Park and Pham [30], which resolved the renowned conjecture of Kahn and Kalai. Let us recall the definition of the expectation threshold and then state the theorem of Park and Pham. Roughly speaking, the expectation threshold $p_{e}(\mathcal{Q})$ is the moment when the expected number of graphs generating $\mathcal{Q}$ hits $1/2$. Formally, $p_{e}(\mathcal{Q})$ is the maximum $p$ such that there exists a set of graphs $\mathcal{Q}^{\prime}$ satisfying $\sum_{G\in\mathcal{Q}^{\prime}}p^{|E(G)|}\leq\frac{1}{2}$ and $\mathcal{Q}\subseteq\langle\mathcal{Q}^{\prime}\rangle$, where $\langle\mathcal{Q}^{\prime}\rangle$ is the upwards closure of $\mathcal{Q}^{\prime}$. Due to Markov’s inequality, $p_{e}(\mathcal{Q})\leq p_{c}(\mathcal{Q})$. Let $n_{\mathcal{Q}}$ be the number of edges in a largest minimal element of $\mathcal{Q}$.
@@
The upper bound in this theorem is tight: there exists an increasing $\mathcal{Q}$ such that $p_{c}(\mathcal{Q})=\Theta(p_{e}(\mathcal{Q})\cdot\log n_{\mathcal{Q}})$. For example, let $\mathcal{Q}$ be the property of being Hamiltonian. It is a routine to check that $p_{e}=\Theta(1/n)$, while $p_{c}=(1+o(1))\frac{\ln n}{n}$ [1, 4, 22, 23, 32]. It is fair to say that the main remaining challenge is to classify properties based on the value of the ratio $\frac{p_{c}(\mathcal{Q})}{p_{e}(\mathcal{Q})}\geq 1$, see, e.g., [3]. In particular, for which properties $\mathcal{Q}$ is $p_{c}(\mathcal{Q})=(1+o(1))p_{e}(\mathcal{Q})$? In this paper, we explore this question for the property of containing a given (unlabelled) regular spanning subgraph.
@@
Let $d\geq 3$ be a fixed constant. Let $F=F(n)$, $n\in\mathbb{N}$, be a sequence of $d$-regular $n$-vertex graphs. We assume that $dn$ is even, let $V(F(n))=\{1,\ldots,n\}$, and use the shorthand notation $[n]:=\{1,\ldots,n\}$. For $\tilde{F}\subset F$, we call a vertex $v\in V(\tilde{F})$ boundary, if it has degree strictly less than $d$ in $\tilde{F}$. We call the vertex boundary of $\tilde{F}$ the set $\partial_{v}(\tilde{F})$ of all boundary vertices and the edge boundary the set $\partial_{e}(\tilde{F})$ of all edges of $F$ between $\partial_{v}(\tilde{F})$ and $V(F)\setminus V(\tilde{F})$. Let the increasing property $\mathcal{Q}_{F}$ be generated by the family $\mathcal{F}_{n}$ of all isomorphic copies of $F$ on $[n]$. For simplicity, we denote $p_{c}(F):=p_{c}(\mathcal{Q}_{F})$ and $p_{e}(F):=p_{e}(\mathcal{Q}_{F})$. We also let $X_{F}$ be the number of graphs from $\mathcal{F}_{n}$ that are subgraphs of $\mathbf{G}\sim G(n,p)$. Since
@@
we get that $p_{c}(F)\geq p_{e}(F)\geq(1-o(1))(e/n)^{2/d}$. On the other hand, if every proper subgraph of $F$ has edge boundary of size at least $d$, then $p_{e}(F)=O(n^{-2/d})$ since it is not hard to see that $\mathcal{F}_{n}$ is $O(n^{-2/d})$-spread, i.e. for every graph $H\subset F$,
@@
if $F$ satisfies a stronger condition on the edge boundary: every $\tilde{F}\subset F$ with $3\leq|V(\tilde{F})|=o(\log n)$ satisfies $|\partial_{e}(\tilde{F})|\geq 2d$. In particular, for powers of a Hamilton cycle [^3] For a graph $G$, its $k$-th power $G^{k}$ is obtained by adding to $G$ edges between vertices that are at distance at most $k$ in the graph metric induced by $G$., this result implies the following: for every $k\geq 3$, the threshold probability for containing the $k$-th power of a Hamilton cycle equals $\Theta(n^{-1/k})$. However, the proof of Riordan does not work for $k=2$ since the minimum size of the edge boundary of a non-trivial subgraph of the square of a Hamilton cycle equals $d+2=6<2d$. We also note that, for a $d$-regular graph to satisfy the condition of Riordan, $d$ should be at least 4. In [25], Kühn and Osthus proved that $n^{-1/2+o(1)}$ is the threshold probability for containing the square of a Hamilton cycle and conjectured that the threshold is actually $\Theta(n^{-1/2})$. In [29], Nenadov and Škorić proved the upper bound $n^{-1/2}(\log n)^{4}$, which was improved to $n^{-1/2}(\log n)^{3}$ by Fischer, Škorić, Steger, and Trujić in [12], and to $n^{-1/2}(\log n)^{2}$ in an unpublished work of Montgomery (see [17]). From Theorem 1.1 (as well as from its weaker fractional version [13]), it follows that the threshold is $O(n^{-1/2}\log n)$. The conjecture of Kühn and Osthus was eventually resolved by Kahn, Narayanan, and Park in 2020 [20]. They further conjectured that $p_{c}(F)=(1+o(1))\sqrt{e/n}$. This paper resolves the conjecture, see Section 1.4.
On the other hand, if $F$ has many small subgraphs with smaller edge boundaries, then it may happen that $p_{c}(F)\gg p_{e}(F)$. For instance, assume that there exists $v=v(n)=o(\log n)$ and a graph $H=H(n)$ on $[v]$ with $\frac{dv}{2}-\frac{d}{2}$ edges such that every vertex of $F$ belongs to a subgraph which is isomorphic to $H$ (clearly, its edge boundary equals $d$). Then, an increasing (at most logarithmic) factor arises — for constant $v$ see the proof in [34] which can be directly generalised to all $v=o(\log n)$. Actually, if we allow even smaller edge boundaries, then the threshold probability may increase by a power of $n$, as well as the expectation threshold. For instance, for every $d$-regular strictly 1-balanced graph $H$ of constant size $v$, the threshold probability for containing an $H$-factor equals $\Theta\left(n^{-2/d+2/(vd)}(\ln n)^{2/(dv)}\right)$ [7, 19].
@@
As follows from the above discussion, our bound on edge boundaries is tight, that is, in some sense, the condition on the edge boundary is the only obstacle in getting the equality (1). Indeed, for every $3\leq v=o(\log n)$ and every $0\leq\ell\leq d$ (such that $dv-\ell$ is even), there exists a graph $F$ that have a subgraph on $v$ vertices with edge boundary of size $\ell$ such that (1) does not hold. The proof of Theorem 1.2 uses similar techniques as in [8, 20] and does not require accurate spread approximations, for this reason we relegate it to Appendix A.
@@
Let $\mathbf{G}\sim G(n,p)$. The threshold probability $p_{c}(\mathcal{Q})$ is called sharp, if, for every constant $\varepsilon>0$, the following is true: if $p>(1+\varepsilon)p_{c}(\mathcal{Q})$, then whp $\mathbf{G}\in\mathcal{Q}$; if $p<(1-\varepsilon)p_{c}(\mathcal{Q})$, then whp $\mathbf{G}\notin\mathcal{Q}$. Friedgut proved [15] that all increasing and isomorphism-closed [^4] In the appendix to this paper, written by Bourgain, an alternative version of this result — one that does not require the symmetry condition — is proved. properties that do not have sharp thresholds are essentially local: with large probability, they are equivalent to the presence of certain property-specific subgraphs of bounded sizes. Omitting technical details, it means that, for some $p=\Theta(p_{c}(\mathcal{Q}))$, $\mathbf{G}\sim G(n,p)$ sprinkled with $\mathbf{G}^{\prime}\sim G(n,\varepsilon p)$, for some $\varepsilon>0$, is less likely to have the property $\mathcal{Q}$, than the union of $\mathbf{G}$ with an independently sampled random clique of some bounded size (see, e.g., [14, Theorem 2.3]). With this remarkable result, it is fairly straightforward to check that the threshold probabilities in Theorem 1.2 are sharp. To the best of our knowledge, there are no general results describing asymptotics of sharp $p_{c}$ for a wide class of increasing properties (say, for $\mathcal{Q}_{F}$, where $F$ are $d$-regular). In this paper, we find asymptotics of $p_{c}(F)$ for most $d$-regular graphs $F$.
@@
Note that, for every $d\geq 4$, almost all $d$-regular graphs satisfy the first condition in Theorem 2 (see [5, 21, 26]). Indeed, for every $d\geq 3$, letting $\mathbf{G}_{d}$ be a random $d$-regular graph on $[n]$, whp $\mathbf{G}_{d}$ is asymmetric [26], implying 1.(a). Moreover, whp $\mathbf{G}_{d}$ has Cheeger constant $\Theta(1)$ [5] and, whenever $d\geq 4$, for every fixed graph $H$ with at least $(d|V(H)|-d)/2$ edges, whp $\mathbf{G}_{d}$ does not have a subgraph isomorphic to $H$ [21], implying 1.(b). If $d=3$, then whp there are no subgraphs with $33\leq v\leq n-$ 3 vertices and the edge boundary of size at most $d$ other than $K_{3}$ and their vertex-complements. Therefore, whp $\mathbf{G}_{d}$ satisfies the first condition, if it does not contain triangles. Also, $k$-th powers of Hamilton cycles for every $k\geq 3$ and toroidal square grids $T_{m,n/m}\cong C_{m}\square C_{n/m}$ for $m\geq 4$ satisfy the second condition as well. It is also easy to see that the second condition implies sharp thresholds for square and triangular lattices (by considering their regular completions). It is worth recalling that thresholds for lattices were investigated in the past: the question of which random graphs contain spanning lattices was raised by Venkatesan and Levin [38]. Thresholds for lattices, up to a polylog-factor, were proved in [11] and, up to a constant factor, were proved in [33] (they also follow from Theorem 1.2 and [8, Corollary 1.5]). Summing up, we get the following corollary that demonstrates that the set of regular graphs $F$ satisfying the restrictions in Theorem 2 is quite broad.
@@
Examples of graphs that do not satisfy restrictions from Theorem 2 include the square of a cycle, toroidal grid $T_{3,n/3}$, and, so-called, 2-overlapping 4-cycles. The latter graph, denoted by $C^{e}_{4,n}$, is obtained from $n/2$ cyclically order copies of $C_{4}$, where two consecutive $C_{4}$ overlap in exactly one edge, whereby each $C_{4}$ overlaps with two copies of $C_{4}$ in opposite edges. Recall that Kahn, Narayanan, and Park in [20] proved (1) for the square of a cycle $F$, resolving the conjecture of Kühn and Osthus [25]. Moreover, using the same method, Espuny Díaz and Person in [10] proved (1) for $F=C^{e}_{4,n}$ answering the question of Frieze [16] (we note that (1) for these graphs also follow both from Theorem 1.2 and [8, Corollary 1.5]).
1.4. Squares of Hamilton cycles.
@@
In 2020, Kahn, Narayanan, and Park [20] conjectured that (2) holds when $F$ is the square of a cycle. Due to the natural barrier in the application of the fragmentation technique that was used to prove (1) for the square of a Hamilton cycle [20], to resolve the conjecture of Kahn and Kalai [30], and to prove the so called spread lemma [2] (see details in Section 1.5), this conjecture attracted significant interest and was reiterated in [17, 31]. In this paper, we prove the conjecture.
@@
Let $F=F(n)$ be the square of an $n$-cycle on $[n]$. Then $F$ satisfies (2).
@@
Let $d\geq 3$ be a constant, $r=r(n)=o(\log\log n/\log\log\log n)$ [^5] We did not try to optimise the bound on $r$ and we believe that it can be improved using our methods, but not beyond $\log n$., and $F=F(n)$ be a sequence of graphs on $[n]$ satisfying the following.
@@
In particular, (2) holds for the two other ‘naive’ cases that Theorem 2 does not cover — for $F=C^{e}_{4,n}$ and $F=T_{3,n/3}$. Indeed, for these two graphs the requirements of Theorem 1.8 are satisfied with $r=4$ and $r=3$, respectively.
@@
In order to prove Theorem 1.7, we introduce a novel technique of choosing fragments — we show that the “cyclic” structure of squares of Hamilton cycles allows to choose rare fragments that are sparser than typical ones, from sufficiently many copies of $F$. The key idea from [20] that proves (1) for squares of cycles is as follows. Fix any $F\in\mathcal{F}_{n}$, before the edges of the uniformly random graph $\mathbf{W}\sim G(n,m)$ are exposed, where $m=\lfloor p{n\choose 2}\rfloor$. Consider $F\cup\mathbf{W}$ and show that whp almost every $F^{\prime}\subset F\cup\mathbf{W}$ from $\mathcal{F}_{n}$ is such that $|F^{\prime}\setminus\mathbf{W}|=O(\sqrt{n})$. It means that we may replace almost every $F\in\mathcal{F}_{n}$ with its fragment $F^{\prime}\setminus\mathbf{W}\subset F$ — for typical $F^{\prime}\subset F\cup\mathbf{W}$ — of size $O(\sqrt{n})$ and expose another $\mathbf{\tilde{W}}\sim G(n,\tilde{m})$ independently where $\tilde{m}$ is large enough to cover at least one fragment whp. Then whp $\mathbf{W}\cup\mathbf{\tilde{W}}$ contains a graph from $\mathcal{F}_{n}$.
However, if a fragment contains a subgraph with edge boundary of size $d+2=6$ with $\Omega(\ln n)$ edges — such a subgraph is exactly the square of a path, we call it closed — then $\tilde{m}=\lceil\delta n\sqrt{n}\rceil$ additional uniformly random edges with a certain bounded from zero $\delta>0$ is not enough to cover it. Unfortunately, we suspect that typical fragments do contain closed subgraphs of size $\Omega(\ln n)$. So, this approach does not allow to get a sharp bound on the threshold. This is the main complication that does not allow to apply directly the fragmentation technique to resolve the conjecture of Kahn, Narayanan, and Park. A careful implementation of the technique gives upper bound on the threshold $(\sqrt{e}+1+o(1))n^{-1/2}$, see Remark 1.3.
@@
The proof of Theorem 2 is much more straightforward — typical fragments are good enough to derive (2). Indeed, both conditions exclude closed subgraph of size $\Omega(\log n)$. We prove the first part of Theorem 2 via $1/\delta$ fragmentation steps. The proof of the second part of Theorem 2 requires only two fragmentations. Actually the usual second moment method works in this case — for the uniform model instead of the binomial, as in the paper of Riordan [33]. It is possible to show that $\mathrm{Var}X_{F}=O((\mathbb{E}X_{F})^{2})$ and then to apply the powerful results of Friedgut [14, 15]. In particular, this strategy was used by Narayanan and Schacht to determine sharp thresholds for non-loose Hamiltonian cycles in random hypergraphs [29]. However, we give the proof of the second part of Theorem 2 using fragmentation for the sake of convenience and coherence.
@@
Let us finally say a few words about our proof of Theorem 1.2. Kahn, Narayanan, and Park in [20] noted that the key observation enabling their proof that the threshold for the appearance of the second power of a Hamilton cycle $F$ equals $p_{c}(F)=\Theta(n^{-1/2})$ is that $\mathcal{F}_{n}$ is $(1+o(1))\sqrt{e/n}$-spread. They refined the concept of spreadness by incorporating the number of components in a subgraph. This refinement was later distilled by Espuny Díaz and Person in [10] under the name of superspreadness and was used to extend the result of Kahn, Narayanan, and Park to a broader class of spanning subgraphs in $G(n,p)$. Somewhat surprisingly, this property allows to prove the weaker result [8, Corollary 1.5] but does not allow to prove Theorem 1.2, as it treats contributions from connected components of both bounded and growing sizes in the same manner. It is also worth noting that Spiro [35] proposed another generalisation of spreadness, which, in particular, recovers the results from [10] and served as a key ingredient in the proofs of [8].